log5(x+7)+log5(x-7)=2
Answer by KnightOwlTutor(293)

we know that 5^2=25
therefore x+7+x-7=25
Simplify
2x=25
Divide both sides by 2
x=12.5

Plug into the equation to doublecheck
12.5+7=19.5
12.5-7=5.5
19.5+5.5=25